# Rethinking Layer Removal: Preserving Critical Components with Task-Aware Singular Value Decomposition

### v1

@@ 5 Discussion
5.1 How Taco-SVD Stabilizes Layer Removal and Enhances Model Consistency
Instabilities Caused by Layer Removal
Layer removal can disrupt the internal consistency of a model, leading to performance degradation. In response, the model attempts to compensate for the loss of removed layers, triggering a process we refer to as self-repair (Rushing and Nanda, 2024). While this self-repair mechanism helps the model maintain partial functionality, it is often noisy and incomplete. The process cannot fully recover the removed components, resulting in higher perplexity and reduced cross-task generalization. This phenomenon is task-dependent, with certain tasks being more sensitive to the loss of critical task-relevant transformations.
Taco-SVD as a Stabilizing Solution
Taco-SVD effectively addresses these issues by preserving key task-critical singular value directions, reducing the need for self-repair and stabilizing the model’s internal consistency. Specifically, Taco-SVD:
  • Retains task-critical singular values, minimizing disruption to the residual stream. This helps maintain the original structure and function of the model, reducing perplexity and improving performance across diverse tasks.
  • Mitigates the reliance on self-repair, ensuring that the model does not need to rely on incomplete compensations. By aligning the retained singular values with task-specific objectives, Taco-SVD facilitates smoother adjustments that preserve both task performance and internal consistency.
5.2 Taco-SVD vs LoRA vs LLM-Streamline
Limitations of fine-tuning Pruned LLMs with LoRA
Parameter-efficient fine-tuning methods such as LoRA introduce additional parameters across the remaining layers to compensate for the removed layers. However, training the pruned model with LoRA is a process of redistributing the function of the pruned layers (Chen et al., 2024), these updates often introduce intruder dimensions, which may lead to potential issues such as Spectral Dilution (Shuttleworth et al., 2024): LoRA’s weight updates alter the original singular value spectrum, the new singular vectors added by LoRA are often orthogonal to the original pre-trained singular vectors, which diminishes the contribution of task-agnostic features learned during pre-training.
Limitations of Layer Replacement with Dense Modules
Methods like LLM-Streamline (Chen et al., 2024) train a dense module from scratch to approximate the functionality of redundant layers within LLMs. However, dense modules possess high-dimensional internal structures, which may cause overfitting in scenarios with limited data.
Taco-SVD’s Generalization Advantage via Low-Rank Learning
In contrast, Taco-SVD avoids such issues by preserving the original spectral structure while aligning it with task-specific objectives.
  • Taco-SVD identifies and retains the singular value directions in the redundant layer that are most critical to task performance. It maintains the original pre-trained singular value spectrum, preventing the introduction of orthogonal updates to the original singular vectors in the remaining layers.
  • Previous literature has demonstrated the effectiveness and practicality of low-rank learning in deep networks (Oymak et al., 2019; Kamalakara et al., 2022; Yaras et al., 2024). Therefore, instead of training a dense module from scratch, Taco-SVD fine-tunes only the minimal task-aware subspaces in redundant layers. It can avoid overfitting by constraining the model to learn within a low-rank structure.
@@ 6 Conclusion
We propose Taco-SVD, a task-aware framework for compressing large language models by retaining task-critical singular value directions. Taco-SVD mitigates the disruptions caused by traditional layer removal and avoids the need for computationally expensive fine-tuning compensation by preserving key task-critical transformations. Experimental results show that Taco-SVD achieves lower perplexity, better task performance, and improved cross-task robustness with minimal resource overhead. This work highlights the effectiveness of aligning low-rank decompositions with task-specific objectives for efficient and robust model compression.
@@ 7 Limitations
While Taco-SVD demonstrates strong performance across various tasks and architectures, we acknowledge certain limitations that highlight opportunities for future research:
Dependence on Calibration Data Taco-SVD relies on calibration data for singular value gradient attribution. In our experiments, using as few as 64 calibration samples from the general-purpose WikiText dataset achieved strong performance across multiple tasks. However, while this approach works well in these settings, we have not yet explored whether using domain-specific calibration data could yield better results in specialized domains. Further testing with such data may provide valuable insights into its potential advantages.
Trade-Off Between Compression and Retained Parameters To ensure robustness and task consistency, Taco-SVD retains a small fraction of low-rank parameters in redundant layers, mitigating disruptions caused by pure layer removal. This design choice introduces a slight trade-off in terms of parameter retention, which could limit its suitability for applications requiring extreme compression ratios. Further reduction techniques may complement Taco-SVD in such scenarios.
Scalability to Extremely Large Models While Taco-SVD performs effectively on models like LLaMA2 and LLaMA3, its application to ultra-large-scale models (e.g., hundreds of billions of parameters) could pose additional computational challenges. Future work could focus on optimizing the algorithm for such architectures, enabling its application to next-generation LLMs.
@@
While Taco-SVD is designed to optimize the efficiency and scalability of large language models, we recognize that such technologies could be misused in applications that may perpetuate harmful biases or deploy models in contexts lacking adequate oversight. To mitigate these risks, we advocate for responsible deployment practices, including thorough testing and monitoring for unintended biases.
@@ A.1 The gradient of singular values
For a calibration dataset $D$, the gradient of a singular value $\sigma_{i}$ with respect to the task loss can be interpreted as the projection of the weight gradient matrix $G$ onto the corresponding singular direction, given by:
@@ A.2 Layer-wise Low-Rank Decomposition
Algorithm 1 Task-Aware SVD for LLM Compression
1:Model $M$ with $N$ layers, Calibration Dataset $D$, Retain Ratio $r$
@@
9:Step 2: Task-aware SVD Compression:
@@ A.3 Experimental Setup and Hyperparameters Configuration
To ensure a fair comparison, all experimental setups are consistent across all methods. Below we describe the experimental setup and hyperparameters configuration in detail.
Training-free Setting
We report the results of Taco-SVD and other training-free baselines under the compression ratio of 20% in Table 1. Specifically, we selected the same 256 samples from the WikiText-2 dataset as calibration data.
Training-based Setting
Table 6 provides the detailed configurations of training-based methods under the compression ratio of 20%. All our experiments are conducted on a single NVIDIA A100 GPU with mixed precision enabled.
@@
Table 6: Experimental setup and hyperparameter configurations.
@@
Method Ratio WikiText-2 PTB Openb. ARC_e WinoG. HellaS. ARC_c PIQA MathQA Average
Dense 0% 5.47 20.83 31.8 69.28 67.09 56.7 39.93 78.34 27.67 52.97
LaCo 18% 39.43 93.98 26.2 48.36 59.19 41.93 32.42 68.72 24.02 42.98
SliceGPT 21.50% 8.63 99.25 22 54.34 60.62 37.4 27.99 63.38 23.18 41.27
ASVD 20% 9.7 88.39 27.8 58.41 61.09 46.97 32.42 71.05 25.09 46.12
Taco-SVD 21.60% 16.12 44.09 24 53.59 62.9 42.94 32.76 71.06 22.88 44.30
LaCo 24% 50.33 102.84 25 45.33 60.06 38.3 29.69 65.51 23.99 41.13
SliceGPT 25.10% 9.54 118.17 20.6 50.91 59.63 35.66 26.41 61.62 23.43 39.75
ASVD 25% 53.48 374.45 17 39.35 53.12 32.64 22.78 61.48 22.95 35.62
Taco-SVD 25% 30.2 60.74 21.6 46.59 61.8 40.41 28.84 66.1 23.18 41.22
LaCo 30% 94.5 122.09 22 42.26 59.43 35.7 28.41 62.89 24.32 39.29
SliceGPT 29.10% 10.91 145.24 18.8 47.56 58.64 34.18 25.34 59.58 23.15 38.18
ASVD 30% 1877.63 3340.65 14.4 27.78 50.59 26.37 21.84 51.58 20.88 30.49
Taco-SVD 30% 35.38 88.96 19.4 42.72 61.17 38.43 27.73 63.98 23.18 39.52
LaCo 36% 138.82 164.02 19 38.34 58.96 33.71 27.13 60.55 23.55 37.32
SliceGPT 33.20% 12.8 176.54 17.8 44.23 57.61 32.28 23.55 58.59 22.68 36.68
ASVD 35% 2599.79 6529.72 13.6 26.81 50.83 25.94 23.12 53.59 19.33 30.46
Taco-SVD 35.00% 68.43 125.79 18.8 40.25 60.35 36.77 27.18 61.11 22.63 38.16
LaCo 39% 386.55 289.98 18.4 33.96 56.99 31.34 26.88 58.6 23.45 35.66
SliceGPT 41.40% 19.1 257.66 15.8 35.94 53.43 29.62 20.82 55.98 22.01 33.37
ASVD 40% NaN NaN 13.4 25.76 51.3 26.1 21.67 54.19 20.64 30.44
Taco-SVD 40.10% 80.42 247.55 18 36.83 59.59 33.31 26.62 59.63 21.88 36.55
@@
Table 7: Zero-shot performance of LLaMA2-7B compressed by Taco-SVD and training-free baselines under 20% to 40% compression. Bold values indicate the best performance.
A.4 Validation Loss Curves
We plot the validation loss curves when training pruned Mistral-7B with 20% compression ratio with different compensation methods. As shown in Figure 5, we can observe that Taco-SVD has better initialization, leading to more stable and faster convergence during fine-tuning.
A.5 Results under High Compression Ratio
In this section, we present the results of Taco-SVD compared with the training-free baselines using LLaMA2-7B under high compression ratios, as shown in Table 7. The results demonstrate that our proposed method consistently outperforms the baselines, even at high compression ratios of up to 40%.
A.6 Robustness of Taco-SVD towards different Calibration Dataset
In this section, we provide details of the ablation studies conducted to investigate the impact of calibration datasets and the number of calibration data used for singular value gradient attribution. Specifically, we selected 512 samples from WikiText-2 (Merity et al., 2017) and C4 (Raffel et al., 2020) as calibration data to assess the performance of Taco-SVD when compressing LLaMA3.1-8B-Instruct under 20% compression ratio. Additionally, we selected 64, 128, 256 and 512 samples from WikiText-2 to examine the robustness of Taco-SVD to the change in the number of calibration data. All calibration data were randomly selected from the training splits of the downstream datasets, ensuring no data leakage. As shown in Figure 6, we can observe that Taco-SVD consistently achieves strong performance, indicating that our method is robust to variations in both the calibration dataset and the number of calibration data. Tables 8 and 9 summarize the results of Taco-SVD when compressing LLaMA3.1-8B-Instruct with different calibration datasets (WikiText-2, C4) and varying numbers of calibration data.
Figure 6: Performance of Taco-SVD on LLaMA3.1-8B-Instruct under 20% compression using (a) different calibration datasets (WikiText-2, C4) and (b) varying amounts of calibration data from WikiText-2. Taco-SVD demonstrates limited sensitivity to calibration data changes, with final task performance varying within 4%.
@@ A.6 Robustness of Taco-SVD towards different Calibration Dataset
Calibration Dataset WikiText-2 PTB Openb. ARC_e WinoG. HellaS. ARC_c PIQA MathQA Average

### v2

@@ 5 Discussion
GRASP preserves representation stability.
Layer removal disrupts the internal spectral structure of LLMs, often introducing inconsistencies in the learned representations. These disruptions lead to degradation in perplexity and loss of linguistic coherence, as the model struggles to compensate for missing transformations. GRASP stabilizes the compression process by selectively retaining representation-critical singular values, ensuring that the model’s residual stream remains structurally consistent. Unlike direct layer removal, which forces the model to undergo self-repair (Rushing and Nanda, 2024), GRASP proactively mitigates this instability by preserving the most influential spectral components, thereby reducing the burden on later layers to reconstruct lost information.
GRASP accelerates adaptation via low-rank learning.
Fine-tuning pruned models using parameter-efficient methods such as LoRA often introduces intruder dimensions into the pre-trained singular spectrum (Shuttleworth et al., 2024), leading to spectral dilution that weakens the contribution of pre-trained features. Similarly, replacing layers with dense modules, as in LLM-Streamline (Chen et al., 2024), requires training high-dimensional structures from scratch, which increases the risk of overfitting when data is limited. In contrast, GRASP identifies and fine-tunes the most essential low-rank subspaces within redundant layers, preserving the original spectral alignment while adapting efficiently to new tasks. Prior research has established the effectiveness of low-rank learning in deep networks (Oymak et al., 2019; Kamalakara et al., 2022), and GRASP leverages this principle to optimize task-specific performance while maintaining pre-trained feature integrity.
@@ 6 Conclusion
We present GRASP, a novel framework for compressing large language models by selectively retaining representation-critical singular value directions. By mitigating the disruptions caused by layer removal, GRASP preserves essential transformations necessary for maintaining model stability and generalization. Experimental results demonstrate that GRASP achieves efficient and robust compression, improving perplexity and downstream task performance compared to existing methods while maintaining minimal computational overhead.
@@ 7 Limitations
While GRASP demonstrates strong performance across various tasks and architectures, we acknowledge certain limitations that highlight opportunities for future research:
Dependence on Calibration Data GRASP relies on representation evaluation dataset for singular value gradient attribution. In our experiments, using as few as 64 samples from the WikiText dataset achieved strong performance across multiple tasks. However, while this approach works well in these settings, we have not yet explored whether using domain-specific dataset could yield better results in specialized domains. Further testing with such data may provide valuable insights into its potential advantages.
Trade-Off Between Compression and Retained Parameters To ensure robustness and task consistency, GRASP retains a small fraction of low-rank parameters in redundant layers, mitigating disruptions caused by pure layer removal. This design choice introduces a slight trade-off in terms of parameter retention, which could limit its suitability for applications requiring extreme compression ratios. Further reduction techniques may complement GRASP in such scenarios.
Scalability to Extremely Large Models While GRASP performs effectively on models like LLaMA2 and LLaMA3, its application to ultra-large-scale models (e.g., hundreds of billions of parameters) could pose additional computational challenges. Future work could focus on optimizing the algorithm for such architectures, enabling its application to next-generation LLMs.
@@
While GRASP is designed to optimize the efficiency and scalability of large language models, we recognize that such technologies could be misused in applications that may perpetuate harmful biases or deploy models in contexts lacking adequate oversight. To mitigate these risks, we advocate for responsible deployment practices, including thorough testing and monitoring for unintended biases.
@@ A.1 The gradient of singular values
For a representation evaluation dataset $D$, the gradient of a singular value $\sigma_{i}$ with respect to the task loss can be interpreted as the projection of the weight gradient matrix $G$ onto the corresponding singular direction, given by:
@@ A.2 Layer-wise Low-Rank Decomposition
Algorithm 1 Gradient-Guided SVD for LLM Compression
1:Model $M$ with $N$ layers, Representation Evaluation Dataset $D$, Retain Ratio $r$
@@
9:Step 2: Gradient-guided SVD Compression:
@@
Method Metric Calibration Data Need Post-Training Training Data Training Dataset Size Training Module
ShortGPT Cosine Similarity WikiText-2 No None None None
LaCo Cosine Similarity WikiText-2 Optional Unpublished 1B Full Parameters
SliceGPT PCA
WikiText-2
Alpaca
Optional Alpaca 5k Full Parameters
Shortened LLaMA
Taylor
Perplexity
BookCorpus Optional
SlimPajama
Alpaca
627B
50k
Full Parameters
LoRA-Adapter
LLM-Streamline Cosine Similarity WikiText-2 Yes SlimPajama 30k
Lightweight
Network
GRASP Cosine Similarity WikiText-2 Optional Alpaca 50k
Low-rank
Modules
Table 7: Comparison of pruning-based LLM compression methods, where the metric indicates the criterion used to identify redundant modules. "Optional" refers to methods that can either work without post-training or recover performance through post-training. Shortened LLaMA consists of two training stages: initial continual pre-training on the SlimPajama dataset, followed by LoRA fine-tuning on the Alpaca dataset.
@@ A.3 Experimental Setup and Hyperparameters Configuration
To ensure a fair comparison, all experimental setups are consistent across all methods. In the following, we describe the experimental setup and hyperparameters configuration in detail.
Metric
For the WikiText-2 and PTB datasets, we use perplexity ($\downarrow$) to assess model’s generative capacity. For the seven common sense reasoning datasets, performance is measured by accuracy ($\uparrow$), and for the GSM8K dataset, performance is evaluated using Exact Match Accuracy ($\uparrow$). All common sense reasoning tasks are tested in zero-shot setting using the LM-Evaluation-Harness framework (Gao et al., 2024), while GSM8K is tested in the five-shot setting.
Hyperparameters Configurations
For post-training compensation, all models compressed by different methods are trained on the Alpaca (Taori et al., 2023) dataset for 1 epoch with a batch size of 32. We use AdamW (Loshchilov and Hutter, 2019) as our optimizer and set the learning rate to $3\times 10^{-4}$. All our experiments are conducted on a single A100 GPU with mixed precision enabled. Table 8 provides the detailed configurations of post-training compensation.
@@
Table 8: Experimental setup and hyperparameters configurations.
A.4 Comparison of Concurrent Pruning-based Methods
We provide a detailed comparison of concurrent pruning-based LLM compression methods, and the results are summarized in Table 7.
A.5 Detailed Results on Common Sense Reasoning Benchmarks
In this section, we provide the detailed results for each benchmark. Table 9 summarizes the results of pruning-based methods without post-training compensation on the common sense reasoning benchmark, while Table 10 presents the results of methods with post-training compensation.
@@ A.5 Detailed Results on Common Sense Reasoning Benchmarks
Model Method Ratio Openb. ARC_e WinoG. HellaS. ARC_c PIQA MathQA Average
Mistral-7B-v0.1 Dense 0.0% 0.33 0.81 0.74 0.61 0.50 0.81 0.36 0.59
LaCo 21.1% 0.20 0.35 0.58 0.26 0.25 0.53 0.24 0.34
ShortGPT 21.1% 0.19 0.57 0.68 0.46 0.37 0.71 0.26 0.46
SliceGPT 20.0% 0.19 0.51 0.59 0.35 0.25 0.61 0.23 0.39
GRASP 20.0% 0.21 0.56 0.68 0.43 0.38 0.67 0.26 0.46
LLaMA2-7B Dense 0.0% 0.32 0.69 0.67 0.57 0.40 0.78 0.28 0.53
LaCo 18.1% 0.26 0.48 0.59 0.42 0.32 0.69 0.24 0.43
ShortGPT 21.1% 0.23 0.49 0.63 0.42 0.31 0.68 0.23 0.43
SliceGPT 21.5% 0.22 0.54 0.61 0.37 0.28 0.63 0.23 0.41
GRASP 21.6% 0.24 0.54 0.63 0.43 0.33 0.71 0.23 0.44
LLaMA-2-13B Dense 0.0% 0.32 0.73 0.70 0.60 0.46 0.79 0.30 0.56
LaCo 19.5% 0.28 0.52 0.63 0.43 0.33 0.70 0.25 0.45
ShortGPT 22.1% 0.24 0.50 0.63 0.46 0.33 0.7 0.24 0.44
SliceGPT 20.0% 0.29 0.59 0.65 0.39 0.32 0.64 0.24 0.45
GRASP 20.0% 0.26 0.61 0.66 0.47 0.35 0.73 0.24 0.47
LLaMA3.1-8B-Instruct Dense 0.0% 0.34 0.82 0.74 0.59 0.52 0.80 0.39 0.60
LaCo 19.0% 0.26 0.49 0.65 0.33 0.30 0.65 0.30 0.42
ShortGPT 21.7% 0.21 0.57 0.66 0.42 0.32 0.67 0.26 0.44
SliceGPT 20.0% 0.15 0.43 0.51 0.30 0.23 0.58 0.22 0.35
GRASP 20.0% 0.22 0.60 0.70 0.44 0.37 0.69 0.28 0.47
@@
Table 9: Zero-shot performance of GRASP and pruning-based without post-training baselines under 20% compression ratio. Results are reported on seven reasoning datasets (individual and average accuracy). Bold values indicate the best performance.
@@
Model Method Ratio Openb. ARC_e WinoG. HellaS. ARC_c PIQA MathQA Average
Mistral-7B Dense 0.0% 0.33 0.81 0.74 0.61 0.50 0.81 0.36 0.59
LoRA 25.1% 0.25 0.69 0.71 0.51 0.40 0.72 0.26 0.51
LS-FFN 25.2% 0.22 0.67 0.68 0.50 0.38 0.70 0.27 0.49
LS-Layer 24.1% 0.22 0.69 0.70 0.45 0.39 0.72 0.27 0.49
GRASP 24.3% 0.26 0.69 0.70 0.52 0.41 0.73 0.27 0.51
LLaMA-2-7B Dense 0.0% 0.32 0.69 0.67 0.57 0.40 0.78 0.28 0.53
LoRA 25.0% 0.26 0.58 0.63 0.48 0.33 0.71 0.24 0.46
LS-FFN 25.7% 0.25 0.56 0.63 0.48 0.33 0.71 0.25 0.46
LS-Layer 24.0% 0.25 0.60 0.62 0.50 0.35 0.73 0.24 0.47
GRASP 24.3% 0.28 0.61 0.65 0.50 0.36 0.72 0.25 0.48
LLaMA-2-13B Dense 0.0% 0.32 0.73 0.70 0.60 0.46 0.79 0.30 0.56
LoRA 25.0% 0.27 0.63 0.67 0.53 0.38 0.74 0.25 0.50
LS-FFN 25.4% 0.28 0.61 0.67 0.54 0.38 0.73 0.25 0.50
LS-Layer 24.4% 0.29 0.62 0.67 0.54 0.40 0.74 0.26 0.50
GRASP 24.2% 0.29 0.63 0.68 0.55 0.39 0.74 0.27 0.51
LLaMA3.1-8B-Instruct Dense 0.0% 0.34 0.82 0.74 0.59 0.52 0.80 0.39 0.60
LoRA 22.3% 0.28 0.70 0.69 0.51 0.41 0.73 0.30 0.52
LS-FFN 22.0% 0.26 0.70 0.70 0.49 0.41 0.72 0.30 0.51
LS-Layer 21.7% 0.27 0.74 0.69 0.51 0.43 0.73 0.32 0.53
GRASP 21.0% 0.28 0.72 0.70 0.52 0.43 0.74 0.31 0.53
Table 10: Performance comparison of GRASP and pruning-based with post-training baselines under 25% compression ratio. LS-FFN and LS-Layer denote two variants of LLM-Streamline. Results are reported on seven reasoning datasets (individual and average accuracy). Bold values indicate the best performance.
A.6 Validation Loss Curves
We plot the validation loss curves when training pruned Mistral-7B with 20% compression ratio with different compensation methods. As shown in Figure 4, we can observe that GRASP has better initialization, leading to more stable and faster convergence during training.
A.7 Evaluation Results on LongBench
In this section, we present the detailed results of LLaMA3.1-8B-Instruct and its compressed version by GRASP under 20% compression ratio on LongBench, which are presented in Table 13. The results illustrate that GRASP with post-training compensation still maintains superior performance on long-form reasoning and complex generative tasks.
A.8 Compression Costs of GRASP
To assess the compressed costs of GRASP, we measure its compression time and peak memory consumption when compressing LLaMA2-7B on an NVIDIA A100 GPU under 25% compression ratio. These metrics, summarized in Table 11, indicate that GRASP achieves a good balance between compression time and memory consumption.
Model
Average
Compression Time(h)
Compression Memory
Consumption(MiB)
LaCO 0.05 45046
ASVD 5.5 17632
SliceGPT 0.6 18195
GRASP 0.16 29050
Table 11: Compression time and peak compression memory usage of GRASP and baselines for LLaMA2-7B under 25% compression ratio on a single A100 GPU.
A.9 Robustness of GRASP towards different Representation Evaluation Dataset
In this section, we provide details of the ablation studies conducted to investigate the impact of representation evaluation datasets and the amount of data used for singular value gradient attribution. Specifically, we selected 512 samples from WikiText-2 (Merity et al., 2017) and C4 (Raffel et al., 2020) as calibration data to assess the performance of GRASP when compressing LLaMA3.1-8B-Instruct under 20% compression ratio. Additionally, we selected 64, 128, 256 and 512 samples from WikiText-2 to examine the robustness of GRASP to the change in the number of calibration data. All calibration data were randomly selected from the training splits of the downstream datasets, ensuring no data leakage. As shown in Figure 5, we can observe that GRASP consistently achieves strong performance, indicating that our method is robust to variations in both the representation evaluation dataset and the number of data. Tables 14 and 15 summarize the results of GRASP when compressing LLaMA3.1-8B-Instruct with different representation evaluation datasets (WikiText-2, C4) and varying numbers of calibration data.
A.10 Robustness of GRASP towards Different Pruning Strategy
In this section, we present the results of ablation experiments conducted with different pruning strategies in GRASP, with the results summarized in Table 12
Pruning Strategy WikiText-2 PTB
Average
Accuracy
Compression
Time(h)
One-shot Pruning 37.86 63.97 47.12 0.16
Iterative Pruning 38.39 72.18 47.13 0.22
Model Summarization Few-shot Learning Synthetic Task
1-1 1-2 1-3 1-4 1-1 2-1 2-2 2-3 2-4 3-1 3-2
LLaMA3.1-8B-Instruct 28.35 20.04 25.85 14.51 63 51.3 39.29 16.5 1.64 9.67 17.17
GRASP 25.76 19.41 25.97 8.99 59.5 67.44 38.41 18 1 10 18
Model One-Doc QA Multi-Doc QA Code Completion Average
4-1 4-2 4-3 4-4 5-1 5-2 5-3 5-4 6-1 6-2 ALL
LLaMA3.1-8B-Instruct 9.96 6.04 18.75 14.27 11.27 11.57 7.65 34.16 56.47 51.34 26.09
GRASP 15.32 20.47 34.52 16.46 31.33 26.18 10.28 29.86 33.74 38.09 26.26
Table 13: Performance comparison of LLaMA3.1-8B-Instruct and its compressed version by GRASP under 20% compression ratio on LongBench. The datasets are grouped as follows: (1-1 to 1-4) denote GovReport, QMSum, MultiNews, and VCSUM; (2-1 to 2-4) denote TREC, TriviaQA, SAMSum, and LSHT; (3-1 to 3-3) denote PassageCount, PassageRetrieval-en, and PassageRetrieval-zh; (4-1 to 4-4) denote NarrativeQA, Qasper, MultiFieldQA-en, and MultiFieldQA-zh; (5-1 to 5-4) denote HotpotQA, 2WikiMultihopQA, MuSiQue, and DuReader; (6-1 to 6-2) denote LCC and RepoBench-P.
Figure 5: Performance of GRASP on LLaMA3.1-8B-Instruct under 20% compression using (a) different representation evaluation datasets (WikiText-2, C4) and (b) varying amounts of calibration data from WikiText-2. GRASP demonstrates limited sensitivity to calibration data changes, with final task performance varying within 4%.
Representation Evaluation Dataset WikiText-2 PTB Openb. ARC_e WinoG. HellaS. ARC_c PIQA MathQA Average